# On toric foliated pairs

### v1

@@ 1. Introduction
This paper is a continuation of [FjS2] and is obviously a generalization of [Fj1]. Throughout this paper, we will work over $\mathbb{C}$, the field of complex numbers. The following theorem is a log canonical generalization of [FjS2, Theorem 1.3] or is a generalization of [Fj1, Theorem 0.1] for toric foliated pairs. Note that our approach in this paper is based on the toric Mori theory (see [R], [M, Chapter 14], [Fj1], and [FjS1]).
@@
If ${\operatorname{rank}}\mathscr{F}=\dim X$ in Theorem 1.4, then it is nothing but the original version of Fujita’s freeness for smooth projective toric varieties. It is natural to formulate Fujita’s freeness conjecture for foliations.
Conjecture 1.5 (Fujita’s freeness conjecture for foliations).
Let $X$ be a smooth projective variety and let $\mathscr{F}$ be a foliation on $X$ with ${\operatorname{rank}}\mathscr{F}=r$. Let $A$ be an ample Cartier divisor on $X$. Then $|K_{\mathscr{F}}+(r+1)A|$ is basepoint-free.
Since any ample Cartier divisor on a smooth projective toric variety is very ample, we have the following statement on Fujita’s very ampleness.
@@ Acknowledgments.
The first author was partially supported by JSPS KAKENHI Grant Numbers JP20H00111, JP21H04994, JP23K20787. The second author was partially supported by JSPS KAKENHI Grant Number JP24K06679.
@@ 2. Preliminaries
A foliation on a normal algebraic variety $X$ is a nonzero saturated subsheaf $\mathscr{F}\subset\mathscr{T}_{X}$ that is closed under the Lie bracket, where $\mathscr{T}_{X}$ is the tangent sheaf of $X$. We note that the rank of the foliation $\mathscr{F}$ means the rank of the coherent sheaf $\mathscr{F}$.
@@
For the details, see [Pn], [W] and [CC]. Although we do not need it in this paper, we make a remark for the reader’s convenience.
Remark 2.3 (see [CC, Proposition 1.12 (2)]).
@@
Let us recall the definition of log canonical toric foliated pairs. We note that if ${\operatorname{rank}}\mathscr{F}=\dim X$ in Definition 2.4 then it is nothing but the usual definition of toric log canonical pairs.
Definition 2.4 (Log canonical toric foliated pairs, [CC, Definition 2.1]).
A toric foliated pair $(\mathscr{F},\Delta)$ on a toric variety $X$ consists of a toric foliation $\mathscr{F}$ and an effective torus invariant $\mathbb{R}$-divisor $\Delta$ on $X$ such that $K_{\mathscr{F}}+\Delta$ is $\mathbb{R}$-Cartier. Let $\pi\colon\widetilde{X}\to X$ be a proper toric birational morphism of toric varieties. Then we can write
@@
where $\widetilde{\mathscr{F}}$ is the induced foliation on $\widetilde{X}$ and the sum is over all $\pi$-exceptional divisors $E$. We call $a(E,\mathscr{F},\Delta)$ the discrepancy of $E$ with respect to $(\mathscr{F},\Delta)$. We put $\iota(E)=0$ if $E$ is $\widetilde{\mathscr{F}}$-invariant and $\iota(E)=1$ otherwise. We say that the pair $(\mathscr{F},\Delta)$ is log canonical if $a(E,\mathscr{F},\Delta)\geq-\iota(E)$ for any proper birational toric morphism $\pi\colon\widetilde{X}\to X$ and for any $\pi$-exceptional prime divisor $E$ on $\widetilde{X}$.
@@
Lemma 2.5 ([CC, Proposition 3.8 (1)]).
A toric foliated pair $(\mathscr{F},\Delta)$ is log canonical if and only if ${\operatorname{Supp}}\Delta\subset{\operatorname{Supp}}K_{\mathscr{F}}$ and the coefficients of $\Delta$ are in $[0,1]$.
@@
By definition, we can easily see that $(\mathscr{F},\Delta)$ is log canonical if and only if
@@
is log canonical in the usual sense. Thus the pair $(\mathscr{F},\Delta)$ is log canonical if and only if ${\operatorname{Supp}}\Delta\subset{\operatorname{Supp}}K_{\mathscr{F}}$ and the coefficients of $\Delta$ are in $[0,1]$. ∎
We close this section with a remark on the minimal model program.
@@ 3. Lemmas on projective bundles
since $(\mathscr{F},\Delta)$ is log canonical (see Lemma 2.5). We assume that $l_{(\mathscr{F},\Delta)}(R)>r$ holds. From now, we will only explain how to modify the proof of [FjS2, Theorem 1.3]. Hence we will freely use the same notation as in the proof of [FjS2, Theorem 1.3]. We put $b_{\rho}=1$ for $\rho\not\subset V$ and $b_{i}:=b_{\rho_{i}}$ with $\rho_{i}:=\mathbb{R}_{\geq 0}v_{i}$ for every $i$. By changing the order, we may assume that
@@
We take a small projective $\mathbb{Q}$-factorialization $\psi\colon X^{\prime}\to X$ (see [Fj1, Corollary 5.9]). Let $\Delta^{\prime}$ be the strict transform of $\Delta$ and let $\mathscr{F}^{\prime}$ be the induced foliation on $X^{\prime}$. By construction, we have $K_{\mathscr{F}^{\prime}}+\Delta^{\prime}=\psi^{*}(K_{\mathscr{F}}+\Delta)$. Let $\varphi_{R}\colon X\to Y$ be the contraction morphism associated to $R$. By considering $\varphi_{R}\circ\psi\colon X^{\prime}\to Y$, we can find an extremal ray $R^{\prime}$ of ${\operatorname{NE}}(X^{\prime})$ such that $\psi_{*}R^{\prime}=R$ and $l_{(\mathscr{F}^{\prime},\Delta^{\prime})}(R^{\prime})>r$. Since $X^{\prime}$ is $\mathbb{Q}$-factorial, the associated contraction $\varphi_{R^{\prime}}\colon X^{\prime}\to Y^{\prime}$ is a $\mathbb{P}^{r}$-bundle, $\mathscr{F}^{\prime}$ is the relative tangent sheaf $\mathscr{T}_{X^{\prime}/Y^{\prime}}$, and the sum of the coefficients of $\Delta^{\prime}$ is less than one by Step 1. Note that we can write
@@
We consider $\psi_{Z^{\prime}}:=\psi|_{Z^{\prime}}\colon Z^{\prime}\to Z:=\psi(Z^{\prime})$. Then any positive-dimensional fiber of $\psi_{Z^{\prime}}$ is rationally chain connected since $\psi_{Z^{\prime}}\colon Z^{\prime}\to Z$ is toric. Let $C$ be a rational curve in a fiber of $\psi_{Z^{\prime}}$. Let $C^{\prime}$ be the normalization of $\varphi_{R^{\prime}}(C)$. By applying Lemma 3.2 to the base change of $\varphi_{R^{\prime}}\colon X^{\prime}\to Y^{\prime}$ by $C^{\prime}\to Y^{\prime}$, we can check that $\varphi^{-1}_{R^{\prime}}(\varphi_{R^{\prime}}(C))\subset E$. Therefore, we obtain
@@
In particular, $\psi_{Z^{\prime}}\colon Z^{\prime}\to Z$ is birational. If $\varphi^{-1}_{R^{\prime}}(\varphi_{R^{\prime}}({\operatorname{Exc}}(\psi_{Z^{\prime}})))\subsetneq E$, then we can take a curve $C$ such that $\psi(C)$ is a point with
@@
Let $C^{\prime}$ be the normalization of $\varphi_{R^{\prime}}(C)$. We apply Lemma 3.3 to the base change of $\varphi_{R^{\prime}}\colon X^{\prime}\to Y^{\prime}$ by $C^{\prime}\to Y^{\prime}$. Then
@@
Hence, we have the desired equality $E=\varphi^{-1}_{R^{\prime}}(\varphi_{R^{\prime}}(E))$. We finish the proof of Claim. ∎
We put $\mathcal{L}_{i,Z^{\prime}}:=(\varphi^{*}_{R^{\prime}}\mathcal{L}_{i})|_{Z^{\prime}}$ for every $i$. Let $C$ be any curve on $Z^{\prime}$ such that $\psi_{Z^{\prime}}(C)$ is a point. Then, by applying Lemma 3.3 as before, we can check that $\mathcal{L}_{i,Z^{\prime}}\cdot C=0$ holds for every $i$. This implies that there exists a line bundle $\mathcal{L}_{i,Z}$ on $Z$ such that $\mathcal{L}_{i,Z^{\prime}}=\psi^{*}_{Z^{\prime}}\mathcal{L}_{i,Z}$ holds for every $i$ since $\psi_{Z^{\prime}}\colon Z^{\prime}\to Z$ is a projective toric birational morphism. Hence, $\mathcal{L}_{i,Z^{\prime}}|_{C}$ is a trivial line bundle for every $i$. Let $C^{\prime}$ be the normalization of $\varphi_{R^{\prime}}(C)$. Then the base change of $\varphi_{R^{\prime}}\colon X^{\prime}\to Y^{\prime}$ by $C^{\prime}\to Y^{\prime}$ is the second projection
(4.14) $\mathbb{P}^{r}\times C^{\prime}=\mathbb{P}_{C^{\prime}}(\mathcal{O}_{C^{\prime}}\oplus\cdots\oplus\mathcal{O}_{C^{\prime}})\to C^{\prime}$
by Lemma 3.3. In particular, we obtain that $\Delta^{\prime}\cdot C^{\dagger}=0$ holds for every curve $C^{\dagger}$ on $X^{\prime}$ such that $\psi(C^{\dagger})$ is a point.
@@
We put $X^{\prime\prime}:=\mathbb{P}_{Y}(\mathcal{M}_{0}\oplus\cdots\oplus\mathcal{M}_{r})$. Then $\varphi_{R^{\prime}}\colon X^{\prime}\to Y^{\prime}$ is the base change of $X^{\prime\prime}\to Y$ by $\psi_{Y^{\prime}}\colon Y^{\prime}\to Y$. We put $\rho\colon X^{\prime}\to X^{\prime\prime}$. Then $K_{X^{\prime}/Y^{\prime}}+\Delta^{\prime}=\rho^{*}(K_{X^{\prime\prime}/Y}+\Delta^{\prime\prime})$ with $\Delta^{\prime\prime}:=\rho_{*}\Delta^{\prime}$. By construction, $K_{X^{\prime}/Y^{\prime}}+\Delta^{\prime}=\psi^{*}(K_{X/Y}+\Delta)$. Note that $X^{\prime\prime}$ and $X$ are isomorphic in codimension one. By construction again, $-(K_{X/Y}+\Delta)$ is ample over $Y$ and $-(K_{X^{\prime\prime}/Y}+\Delta^{\prime\prime})$ is also ample over $Y$. Hence, $X$ is isomorphic to $X^{\prime\prime}$ over $Y$. This is what we wanted.
@@
5. Proofs of Corollary 1.2, Theorems 1.3, 1.4, 1.6, and 1.7

### v3

@@ 1. Introduction
The basics of toric foliations and toric foliated minimal model program were already studied in [Pan], [S, Section 10], [W], [CC], and so on. In [FjS2], we discussed lengths of extremal rational curves for toric foliations on projective $\mathbb{Q}$-factorial toric varieties. This paper is a continuation of [FjS2] and is obviously a generalization of [Fj1]. Throughout this paper, we will work over $\mathbb{C}$, the field of complex numbers. The following theorem is a log canonical generalization of [FjS2, Theorem 1.3] or is a generalization of [Fj1, Theorem 0.1] for toric foliated pairs. Note that our approach in this paper is based on the toric Mori theory (see [R], [M, Chapter 14], [Fj1], and [FjS1]).
@@
If ${\operatorname{rank}}\mathscr{F}=\dim X$ in Theorem 1.4, then it is nothing but the original version of Fujita’s freeness for smooth projective toric varieties. Since any ample Cartier divisor on a smooth projective toric variety is very ample, we have the following statement on Fujita’s very ampleness.
@@ Acknowledgments.
The first author was partially supported by JSPS KAKENHI Grant Numbers JP20H00111, JP21H04994, JP23K20787. The second author was partially supported by JSPS KAKENHI Grant Number JP24K06679. The authors would like to thank Fanjun Meng for pointing out that a conjecture in the original version is obviously wrong. They also would like to thank the referee very much for useful comments.
@@ 2. Preliminaries
A foliation on a normal algebraic variety $X$ is a saturated subsheaf $\mathscr{F}\subset\mathscr{T}_{X}$ that is closed under the Lie bracket, where $\mathscr{T}_{X}$ is the tangent sheaf of $X$. We note that the rank of the foliation $\mathscr{F}$ means the rank of the coherent sheaf $\mathscr{F}$.
@@
In this paper, we are mainly interested in the case where $X$ is not $\mathbb{Q}$-factorial.
Remark 2.3.
Let $\mathscr{F}$ be a toric foliation on a (not necessarily $\mathbb{Q}$-factorial) toric variety $X$. Then we can define $K_{\mathscr{F}}$ as follows. We consider the smooth locus $X_{\mathrm{sm}}$ of $X$. Since $X_{\mathrm{sm}}$ is a smooth toric variety, we can define $K_{\mathscr{F}}$ on $X_{\mathrm{sm}}$ as in Theorem 2.2. Since ${\operatorname{codim}}_{X}(X\setminus X_{\mathrm{sm}})\geq 2$, we can extend it to the whole $X$ and obtain a well-defined torus invariant Weil divisor $K_{\mathscr{F}}$ on $X$.
For the details of toric foliations, see [Pan], [W] and [CC]. We make a remark for the reader’s convenience.
Remark 2.4 (see [CC, Corollary 3.3]).
@@
Let us introduce the notion of torically log canonical toric foliated pairs. We note that if ${\operatorname{rank}}\mathscr{F}=\dim X$ in Definition 2.5 then it is nothing but the usual definition of toric log canonical pairs.
Definition 2.5 (Torically log canonical toric foliated pairs).
A toric foliated pair $(\mathscr{F},\Delta)$ on a toric variety $X$ consists of a toric foliation $\mathscr{F}$ and an effective torus invariant $\mathbb{R}$-divisor $\Delta$ on $X$ such that $K_{\mathscr{F}}+\Delta$ is $\mathbb{R}$-Cartier. Let $\pi\colon\widetilde{X}\to X$ be a proper birational toric morphism of toric varieties. Then we can write
@@
where $\widetilde{\mathscr{F}}$ is the induced foliation on $\widetilde{X}$ and the sum is over all $\pi$-exceptional divisors $E$. We call $a(E,\mathscr{F},\Delta)$ the discrepancy of $E$ with respect to $(\mathscr{F},\Delta)$. We put $\iota(E)=0$ if $E$ is $\widetilde{\mathscr{F}}$-invariant and $\iota(E)=1$ otherwise. We say that the pair $(\mathscr{F},\Delta)$ is torically log canonical if $a(E,\mathscr{F},\Delta)\geq-\iota(E)$ for any proper birational toric morphism $\pi\colon\widetilde{X}\to X$ and for any $\pi$-exceptional prime divisor $E$ on $\widetilde{X}$.
@@
Lemma 2.6.
A toric foliated pair $(\mathscr{F},\Delta)$ is torically log canonical if and only if ${\operatorname{Supp}}\Delta\subset{\operatorname{Supp}}K_{\mathscr{F}}$ and the coefficients of $\Delta$ are in $[0,1]$.
@@
By definition, we can easily see that $(\mathscr{F},\Delta)$ is torically log canonical if and only if
@@
is log canonical in the usual sense. Thus the pair $(\mathscr{F},\Delta)$ is torically log canonical if and only if ${\operatorname{Supp}}\Delta\subset{\operatorname{Supp}}K_{\mathscr{F}}$ and the coefficients of $\Delta$ are in $[0,1]$. ∎
Although it is nontrivial, we see that our definition of torically log canonical toric foliated pairs coincides with the log canonicity of toric foliated pairs in the usual sense. In this paper, Definition 2.5 and Lemma 2.6 are sufficient for our purposes.
Theorem 2.7 ([CC, Proposition 4.31]).
Let $(\mathscr{F},\Delta)$ be a toric foliated pair. Then $(\mathscr{F},\Delta)$ is torically log canonical if and only if $(\mathscr{F},\Delta)$ is log canonical in the usual sense for foliated pairs.
Proof.
This follows from Lemma 2.6 and [CC, Proposition 4.31]. ∎
For the precise definition of log canonical foliated pairs, see, for example, [CC, Definition 4.1]. By Theorem 2.7, we can simply say that a toric foliated pair $(\mathscr{F},\Delta)$ is log canonical when it is torically log canonical. We close this section with a remark on the minimal model program.
@@ 3. Lemmas on projective bundles
since $(\mathscr{F},\Delta)$ is log canonical (see Lemma 2.6 and Theorem 2.7). We assume that $l_{(\mathscr{F},\Delta)}(R)>r$ holds. From now, we will only explain how to modify the proof of [FjS2, Theorem 1.3]. Hence we will freely use the same notation as in the proof of [FjS2, Theorem 1.3]. We put $b_{\rho}=1$ for $\rho\not\subset V$ and $b_{i}:=b_{\rho_{i}}$ with $\rho_{i}:=\mathbb{R}_{\geq 0}v_{i}$ for every $i$. By changing the order, we may assume that
@@
We take a small projective $\mathbb{Q}$-factorialization $\psi\colon X^{\prime}\to X$ (see [Fj1, Corollary 5.9]). Let $\Delta^{\prime}$ be the strict transform of $\Delta$ and let $\mathscr{F}^{\prime}$ be the induced foliation on $X^{\prime}$. By construction, we have $K_{\mathscr{F}^{\prime}}+\Delta^{\prime}=\psi^{*}(K_{\mathscr{F}}+\Delta)$. Let $\varphi_{R}\colon X\to Y$ be the contraction morphism associated to $R$ (see [FjS1, Theorem 4.5]). By considering $\varphi_{R}\circ\psi\colon X^{\prime}\to Y$, we can find an extremal ray $R^{\prime}$ of ${\operatorname{NE}}(X^{\prime})$ such that $\psi_{*}R^{\prime}=R$ and $l_{(\mathscr{F}^{\prime},\Delta^{\prime})}(R^{\prime})>r$. Since $X^{\prime}$ is $\mathbb{Q}$-factorial, the associated contraction $\varphi_{R^{\prime}}\colon X^{\prime}\to Y^{\prime}$ is a $\mathbb{P}^{r}$-bundle, $\mathscr{F}^{\prime}$ is the relative tangent sheaf $\mathscr{T}_{X^{\prime}/Y^{\prime}}$, and the sum of the coefficients of $\Delta^{\prime}$ is less than one by Step 1. Note that we can write
@@
We consider $\psi_{Z^{\prime}}:=\psi|_{Z^{\prime}}\colon Z^{\prime}\to Z:=\psi(Z^{\prime})$. Then any positive-dimensional fiber of $\psi_{Z^{\prime}}$ is rationally chain connected since $\psi_{Z^{\prime}}\colon Z^{\prime}\to Z$ is toric. Let $C$ be a rational curve in a fiber of $\psi_{Z^{\prime}}$. Let $C^{\prime}$ be the normalization of $\varphi_{R^{\prime}}(C)$. We consider the base change of $\varphi_{R^{\prime}}\colon X^{\prime}\to Y^{\prime}$ by $\mathbb{P}^{1}\simeq C^{\prime}\to Y^{\prime}$. Then, by Lemma 3.2, we have the following commutative diagram:
[…]
where $p_{1}$ and $p_{2}$ are projections. Thus, the numerical equivalence class of $\alpha(p^{-1}_{1}(P))$ is independent of $P\in\mathbb{P}^{r}$. This implies that $\alpha(p^{-1}_{1}(P))$ is numerically equivalent to $C$ for every $P\in\mathbb{P}^{r}$. Hence $\psi\colon X^{\prime}\to X$ contracts $\alpha(p^{-1}_{1}(P))$ to a point for every $P\in\mathbb{P}^{r}$. Thus, we obtain $\varphi^{-1}_{R^{\prime}}(\varphi_{R^{\prime}}(C))\subset E$. Therefore, we obtain
@@
If $\psi_{Z}\colon Z^{\prime}\to Z$ is not birational, then ${\operatorname{Exc}}(\psi_{Z^{\prime}})=Z^{\prime}$. This implies $X\subset E$ by (4.10). This is obviously a contradiction. Hence we obtain that $\psi_{Z^{\prime}}\colon Z^{\prime}\to Z$ is birational. If $\varphi^{-1}_{R^{\prime}}(\varphi_{R^{\prime}}({\operatorname{Exc}}(\psi_{Z^{\prime}})))\subsetneq E$, then we can take a curve $C$ such that $\psi(C)$ is a point with
@@
Let $C^{\prime}$ be the normalization of $\varphi_{R^{\prime}}(C)$. We consider the base change of $\varphi_{R^{\prime}}\colon X^{\prime}\to Y^{\prime}$ by $C^{\prime}\to Y^{\prime}$. Then we have the following commutative diagram:
[…]
where $X^{\prime}_{C^{\prime}}\to C^{\prime}$ is the base change of $X^{\prime}\to Y^{\prime}$ by $C^{\prime}\to Y^{\prime}$. By construction, we can take a curve $C^{\flat}$ on $X^{\prime}_{C^{\prime}}$ such that $\beta(C^{\flat})=C$. Since $X^{\prime}_{C^{\prime}}$ is a $\mathbb{P}^{r}$-bundle over $C^{\prime}$, the Picard number of $X^{\prime}_{C^{\prime}}$ is two. Moreover, $X^{\prime}_{C^{\prime}}$ has two nontrivial contractions $X^{\prime}_{C^{\prime}}\to C^{\prime}$ and $\psi\circ\alpha\colon X^{\prime}_{C^{\prime}}\to W$, where $W$ is the normalization of $(\psi\circ\alpha)(X^{\prime}_{C^{\prime}})$. Thus $\overline{{\operatorname{NE}}}(X^{\prime}_{C^{\prime}})$ has an extremal ray $Q$ spanned by $C^{\flat}$ since $C^{\flat}$ is contracted by $\psi\circ\alpha$. Note that $C^{\flat}\cdot(K_{X^{\prime}_{C^{\prime}}/C^{\prime}}+\Delta^{\prime}_{C^{\prime}})=0$, where $K_{X^{\prime}_{C^{\prime}}/C^{\prime}}+\Delta^{\prime}_{C^{\prime}}=\alpha^{*}(K_{X^{\prime}/Y^{\prime}}+\Delta^{\prime})$. By Lemma 3.3, we can take a curve $C^{\sharp}$ on $X^{\prime}_{C^{\prime}}$ such that the numerical equivalence class of $C^{\sharp}$ is in $Q$ and $\beta(C^{\sharp})=\varphi^{-1}_{R^{\prime}}(\varphi_{R^{\prime}}(C))\cap Z^{\prime}$. Thus the curve $\varphi^{-1}_{R^{\prime}}(\varphi_{R^{\prime}}(C))\cap Z^{\prime}$ is contracted by $\psi$, that is, $\varphi^{-1}_{R^{\prime}}(\varphi_{R^{\prime}}(C))\cap Z^{\prime}\subset{\operatorname{Exc}}(\psi_{Z^{\prime}})$. Then
@@
Therefore, we have the desired equality $E=\varphi^{-1}_{R^{\prime}}(\varphi_{R^{\prime}}(E))$. We finish the proof of Claim. ∎
We put $\mathcal{L}_{i,Z^{\prime}}:=(\varphi^{*}_{R^{\prime}}\mathcal{L}_{i})|_{Z^{\prime}}$ for every $i$. Let $C$ be any curve on $Z^{\prime}$ such that $\psi_{Z^{\prime}}(C)$ is a point. Let $C^{\prime}$ be the normalization of $\varphi_{R^{\prime}}(C)$. We consider the commutative diagram (4.12) as before. Then we can check that $\mathcal{L}_{i,Z^{\prime}}\cdot C=0$ for every $i$ by applying Lemma 3.3 to $X^{\prime}_{C^{\prime}}\to C^{\prime}$. This implies that there exists a line bundle $\mathcal{L}_{i,Z}$ on $Z$ such that $\mathcal{L}_{i,Z^{\prime}}=\psi^{*}_{Z^{\prime}}\mathcal{L}_{i,Z}$ holds for every $i$ since $\psi_{Z^{\prime}}\colon Z^{\prime}\to Z$ is a projective birational toric morphism. Hence, $\mathcal{L}_{i,Z^{\prime}}|_{C}$ is a trivial line bundle for every $i$. Then $X^{\prime}_{C^{\prime}}\to C^{\prime}$, which is the base change of $\varphi_{R^{\prime}}\colon X^{\prime}\to Y^{\prime}$ by $C^{\prime}\to Y^{\prime}$, is the second projection
(4.16) $\mathbb{P}^{r}\times C^{\prime}=\mathbb{P}_{C^{\prime}}(\mathcal{O}_{C^{\prime}}\oplus\cdots\oplus\mathcal{O}_{C^{\prime}})\to C^{\prime}.$
In particular, we obtain that $\Delta^{\prime}\cdot C^{\dagger}=0$ holds for every curve $C^{\dagger}$ on $X^{\prime}$ such that $\psi(C^{\dagger})$ is a point.
@@
We put $X^{\prime\prime}:=\mathbb{P}_{Y}(\mathcal{M}_{0}\oplus\cdots\oplus\mathcal{M}_{r})$. Then $\varphi_{R^{\prime}}\colon X^{\prime}\to Y^{\prime}$ is the base change of $X^{\prime\prime}\to Y$ by $\psi_{Y^{\prime}}\colon Y^{\prime}\to Y$. We put $\rho\colon X^{\prime}\to X^{\prime\prime}$. Then $K_{X^{\prime}/Y^{\prime}}+\Delta^{\prime}=\rho^{*}(K_{X^{\prime\prime}/Y}+\Delta^{\prime\prime})$ with $\Delta^{\prime\prime}:=\rho_{*}\Delta^{\prime}$. By construction, $K_{X^{\prime}/Y^{\prime}}+\Delta^{\prime}=\psi^{*}(K_{X/Y}+\Delta)$. We set $B:=(\psi_{Y^{\prime}}\circ\varphi_{R^{\prime}})(E)$. Then $B$ is a closed subset of $Y$ with ${\operatorname{codim}}_{Y}B\geq 2$. By Claim and the construction of $\psi_{Y^{\prime}}$, we have $E=(\psi_{Y^{\prime}}\circ\varphi_{R^{\prime}})^{-1}(B)$. Thus we obtain the following commutative diagram:
[…]
Note that ${\operatorname{codim}}_{X^{\prime}}E\geq 2$, ${\operatorname{codim}}_{Y^{\prime}}\varphi_{R^{\prime}}(E)\geq 2$, and ${\operatorname{codim}}_{X}\psi(E)\geq 2$. By the above diagram, we can easily check that $X^{\prime\prime}$ and $X$ are isomorphic in codimension one. By construction again, $-(K_{X/Y}+\Delta)$ is ample over $Y$ and $-(K_{X^{\prime\prime}/Y}+\Delta^{\prime\prime})$ is also ample over $Y$. Hence, $X$ is isomorphic to $X^{\prime\prime}$ over $Y$. This is what we wanted.
@@
Remark 4.1.
The authors do not know how to prove Theorem 1.1 using only toric geometry when $X$ is not $\mathbb{Q}$-factorial.
@@
5. Proofs of Corollary 1.2, Theorems 1.3, 1.4, 1.5, and 1.6
@@
We learned the following example from Professor Fanjun Meng.
Example 5.1.
For any positive integer $n$, we can construct a smooth projective surface $Y$ and an ample Cartier divisor $H$ on $Y$ such that $|nH|$ is not basepoint-free (see [L, Examples 5.1.18 and 5.2.1]). We take an elliptic curve $E$ and a point $P$ of $E$. We put $X:=Y\times E$ and $A:=p^{*}_{1}H+p^{*}_{2}P$, where $p_{1}\colon X\to Y$ and $p_{2}\colon X\to E$ are projections. Then $A$ is an ample Cartier divisor on $X$ such that $|nA|$ is not basepoint-free. We consider $\pi:=p_{1}\colon X\to Y$ and $\mathscr{F}:=\mathscr{T}_{X/Y}$, that is, $\mathscr{F}$ is the relative tangent sheaf of $\pi\colon X\to Y$. Then the canonical bundle $\mathcal{O}_{X}(K_{\mathscr{F}})$ of $\mathscr{F}$ is trivial. In this case, $|K_{\mathscr{F}}+nA|$ is not basepoint-free. Hence we cannot formulate Fujita’s freeness conjecture for foliations naively.